OpenStudy (anonymous):

James and Peter are active members of the Student Guild and are graduating in a class of 27 students in total. A committee of five members to graduation must be formed to organize the festivities. How many committees can be formed so that James and Peter are members?

OpenStudy (anonymous):

Well Total number of students=27 Number of class students for a party=5 So number of committees that can be formed=27/5 = 5 at the maximum.

OpenStudy (bahrom7893):

No tI got this!

OpenStudy (kinggeorge):

Since both James and Peter have to be chose, that leaves 25 students left, and 3 spots on the committee left. Thus, the answer should be\[\binom{25}{3}\]or 25-choose-3

OpenStudy (bahrom7893):

mehhh king got there first.. i was about to post that.. lol

OpenStudy (anonymous):

.............

OpenStudy (anonymous):

@How many parties can be formed if James and Peter ARE INCLUDED in each group or?

OpenStudy (anonymous):

Wait, so each group requires 5 members, James and Peter are members already so, 27-2=25 members Moreover, each group would then have 5-2=3 members so, it can be the way KingGeorge described it, though I'm not sure about if its right..anyways.

OpenStudy (anonymous):

veng. stop..it..the king won..lol

OpenStudy (anonymous):

He won but that's not stopping me from celebrating his victory, is it? :)

OpenStudy (bahrom7893):

I always win..

OpenStudy (anonymous):

whats the problem, hershey? veng helped me understand

OpenStudy (anonymous):

There are 27 members in all, including James and Peter. So, let first group have 3+James+Peter. I still think that only 5 groups can be formed with James and Peter being in one of them.

OpenStudy (anonymous):

Hey King, by that you mean: 25!/3!22! ? cuz the result is huge

OpenStudy (kinggeorge):

\[\binom{25}{3}= {25! \over {3!}(25-3)!}={25! \over {3! 22!}}\]And that is a very huge number.

OpenStudy (kinggeorge):

Well not too huge actually, only 2300

OpenStudy (anonymous):

@King: They haven't asked how many ways can the commities be formed so you sure that's a viable answer?

OpenStudy (anonymous):

2300 sounds weird

OpenStudy (kinggeorge):

2300 is correct because....(wait for next post)

OpenStudy (kinggeorge):

out of those 25 people left to choose from, let's number them from 1 to 25. We can only choose three of them. That means, we have 25 choices for the first person, 24 for the next and 23 for the last. \[25*24*23 = {25! \over 22!}\]Then, each of these 3 person groups can be ordered 3! different ways, so you have to divide by 3!\[{25! \over 22!} \times {1 \over 3!} = {25! \over 3! 22!} = 2300\]

OpenStudy (anonymous):

Right, that's the formula for finding the number of combinations which can be possible for the groups.. @Mel: How much marks is this question for?

OpenStudy (anonymous):

Yep, I get it now. Thanks, King! mmm, not sure what you mean, ven.

OpenStudy (anonymous):

Nevermind, guess it was the number of combinations possible for the groups..misunderstood the question lol.

OpenStudy (anonymous):

Yea, me too. But that's correct, I have the options here :P check them out: a) 9288 b) 3276 c) 9828 d) 2300

OpenStudy (anonymous):

Yeah well 2300 is already proved to be the answer :P

OpenStudy (anonymous):

Yes, what I meant is that they are all ''huge'' numbers.